For formation of 50 seed how many minimum meiotic divisions are required?

Grade:11

1 Answers

Arun
4 years ago

In a meiotic division, we have two phases: Meiosis I and meiosis II. So, we get a total of four haploid cells from one diploid cell.

Hence, for the formation of one seed:

  • Pollen Grain (Microsporogenesis): Four pollen grains are produced as a result of one meiotic division in a pollen sac.
    • 4 pollen grains= 1 meiotic division
    • 1 pollen grain= 1/4 meiotic division
    • 50 pollen grains= 50/4 meiotic divisions
  • Egg (Megasporogenesis): In the megasporangium, only one functional egg cell is produced, because of the degeneration of three others.
    • 1 egg cell= 1 meiotic division
    • 50 egg cells = 50 meiotic divisions

To get the total number of meiotic divisions, we add both the aforementioned values,

Total number of meiotic divisions= 50/4 +50 = 62.5 meiotic divisions.